OpenStudy (pokemon23):

How do I solve this 5x^-4

OpenStudy (anonymous):

\[\frac{5}{x^4}\]

OpenStudy (mertsj):

There's nothing to solve

OpenStudy (anonymous):

what moneybird said :)

OpenStudy (pokemon23):

simplify each expression?

OpenStudy (mertsj):

How do I solve this? 12

OpenStudy (anonymous):

no, u need to get rid of the neg. exponent

OpenStudy (pokemon23):

wouldn't it be 1/5x^4

OpenStudy (anonymous):

Is it (5x)^-4 or 5x^-4

OpenStudy (pokemon23):

5x^-4

OpenStudy (anonymous):

then it would not be 1/5x^4

OpenStudy (anonymous):

lol. how do i solve \[2x\]

OpenStudy (pokemon23):

so I'm correct :)

OpenStudy (pokemon23):

satellite not funny :'(

OpenStudy (akshay_budhkar):

the 5 doesnt go to the denominator. the constant doesn't.. only the power to which the variable is raised gets changed \[k(x)^-n=k/x^n\]

OpenStudy (akshay_budhkar):

lol satellite haha

OpenStudy (anonymous):

no but the point is clear right? you can solve an equation. one does not "solve" an expression. you can "write using positive exponents" or "multiply" or "combine like terms" but you cannot "solve"

OpenStudy (pokemon23):

how is it 5/x^4

OpenStudy (pokemon23):

yes I get it satellite73

OpenStudy (anonymous):

i.e you can say \[5x^{-4}=\frac{5}{x^4}\] but that is just rewriting, it is not "solving"

OpenStudy (anonymous):

reason the "5" staying in the numerator is because the exponent is only applied to the"x" not "5x" if it was \[(5x)^{-4}\] then you would get \[\frac{1}{(5x)^4}\]